when x grams of carbon are heated with y grams of oxygen in a closed vessel, no solid residue is leftbehind. which one of the following statements is correct? (a) must lie between 1.33 and 2.67 (b) must be greater than 2.67 (c) must be less than 1.33 (d) must be greater than 1.33

Grade:11

1 Answers

Dheeraj
5 years ago
Ans is d
 
Since in a closed container and in limited oxygen
Following reaction will take place
C+O2=CO
Using POAC
 On carbon atom
x/12=1*mole of CO________(1)
 
On oxygen atom
2*y/32=1*mole of CO____(2)_
From 1 and 2
 
x/12=y/16
y/x=4/3=1.33
Thankyou